A car has an initial velocity of 30 m/s and an acceleration of -5 m/s2. Find its displacement when it comes to stop.


1
Expert's answer
2021-09-27T09:02:03-0400

We use the information provided and we find first the time (for the car to stop, this occurs when Vf = 0m/s) and then the displacement after substitution for a movement with linear acceleration:


"d_0=0\,m
\\ a=-5\frac{m}{s^2}
\\ v_i=30\frac{m}{s}
\\ v_f=0\frac{m}{s}
\\v_f=v_i+at 
\\ \implies t=\cfrac{v_f-v_i}{a}=\cfrac{(0-30)\frac{m}{s}}{-5\frac{m}{s^2}}=6\,s
\\ \text{Now we substitute t=6s and we find:}
\\d=d_0+v_0t+\cfrac{at^2}{2}
\\d=0\,m+(30\frac{m}{s})(6\,s)+\cfrac{-5\frac{m}{s^2}(6\,s)^2}{2}
\\ \implies d=0\,m+180\,m-90\,m=90\,m"


After substitution, we find that the displacement when the vehicle comes to stop is about 90 m.



Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.